# Spurious-Aware Prototype Refinement for Reliable Out-of-Distribution Detection

### v1

@@ Appendix C Performance on Non-Spurious OOD Datasets
GNorm $90.7_{{\color[rgb]{.5,.5,.5}\pm 0.6}}$ $70.8_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $64.1_{{\color[rgb]{.5,.5,.5}\pm 2.1}}$ $97.3_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $99.0_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $84.4$
ReAct $86.8_{{\color[rgb]{.5,.5,.5}\pm 5.2}}$ $66.1_{{\color[rgb]{.5,.5,.5}\pm 7.5}}$ $50.0_{{\color[rgb]{.5,.5,.5}\pm 8.7}}$ $96.3_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $99.0_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $79.6$
VIM $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $99.6_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $99.9_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $99.6_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $99.8$
MDS $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $99.4_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $97.5_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $99.4$
@@
SHE $96.4_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $98.6_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $99.1_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $94.3_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $98.5_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $97.4$
@@
GNorm $97.2_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $49.7_{{\color[rgb]{.5,.5,.5}\pm 1.1}}$ $59.6_{{\color[rgb]{.5,.5,.5}\pm 2.0}}$ $98.4_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $99.0_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $80.8$
ReAct $96.1_{{\color[rgb]{.5,.5,.5}\pm 1.5}}$ $53.9_{{\color[rgb]{.5,.5,.5}\pm 4.6}}$ $40.6_{{\color[rgb]{.5,.5,.5}\pm 5.9}}$ $98.0_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $99.0_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $77.5$
VIM $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $99.2_{{\color[rgb]{.5,.5,.5}\pm 0.7}}$ $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $99.7_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $99.8$
MDS $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $99.7_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $98.0_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $99.5$
@@
ReAct $57.9_{{\color[rgb]{.5,.5,.5}\pm 15.9}}$ $78.6_{{\color[rgb]{.5,.5,.5}\pm 7.4}}$ $65.0_{{\color[rgb]{.5,.5,.5}\pm 8.0}}$ $91.8_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $99.0_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $78.5$
VIM $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $99.8_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $99.9_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $99.2_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $99.8$
MDS $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $98.8_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $96.5_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $99.1$
@@
SHE $91.6_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $99.4_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $99.5_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $97.0_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $98.3_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $97.2$
@@
We evaluate the performance of post-hoc methods in the non-spurious OOD (NSP-OOD) detection setting. For this evaluation, the OOD samples are drawn from a collection of standard benchmark datasets: SVHN [63], LSUN (resized)[64], and iSUN[65]. The in-distribution (ID) context for each evaluation varies across the five datasets used in our primary spurious correlation experiments: Waterbirds (WB), CelebA (CA), UrbanCars (UC), Animals MetaCoCo (AMC), and Spurious ImageNet (SpI). Specifically, for each of these five ID contexts, the pretrained ResNet-50 backbone is employed, and the post-hoc OOD detection methods are subsequently set up using the features derived from this backbone.
The performance metrics reported in Tables 8 and 9 are the average performance when distinguishing each ID dataset from the NSP-OOD samples, averaged across SVHN, LSUN, and iSUN. This setup assesses the general OOD detection capability of the post-hoc methods when the primary challenge is not spurious correlations shared between ID and OOD, but the ID datasets still contain inherent biases. As observed in the tables, distance-based methods such as SPROD and KNN consistently achieve near-perfect scores across all metrics (AUROC, FPR@95, AUPR-In, and AUPR-Out) and ID contexts. Other methods like VIM and MDS also show strong performance.
Appendix D Ablation Study on SPROD Stages
@@
Appendix E Backbone Experiments
@@
We analyze the performance of each method using four complementary metrics (AUROC, FPR@95, AUPR-IN, and AUPR-OUT) to evaluate different aspects of OOD detection performance. Tables 10, 11, 12, and 13 summarize the performance of our post-hoc method across various backbone architectures, averaged over the five SP-OOD datasets studied in this work. Detailed results for each individual backbone across the different datasets are presented in the subsequent tables.
@@
Overall, we observe that SPROD consistently achieves strong performance across all backbone architectures, often with a notable margin. KNN emerges as the second-best method, suggesting that simpler approaches can be counterintuitively effective; however, its performance is highly sensitive to the choice of hyperparameters. MDS is also among the top-performing methods; it is a metric-based approach similar to SPROD but employs a more complex model by estimating class-specific covariance matrices. This added complexity, while potentially beneficial, may increase the risk of overfitting, especially in OOD settings. In addition, output-based methods show notable drops in performance on several benchmarks, highlighting their susceptibility to distributional shifts. Among the evaluated backbones, ConvNeXt-B and Swin-B stand out as frozen feature extractors with superior performance for post-hoc OOD detection.
@@
MSP $68.1_{{\color[rgb]{.5,.5,.5}\pm 0.7}}$ $47.6_{{\color[rgb]{.5,.5,.5}\pm 2.9}}$ $40.2_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $72.4_{{\color[rgb]{.5,.5,.5}\pm 0.6}}$ $77.3_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $61.1$
@@
ReAct $75.9_{{\color[rgb]{.5,.5,.5}\pm 4.1}}$ $40.8_{{\color[rgb]{.5,.5,.5}\pm 3.7}}$ $44.0_{{\color[rgb]{.5,.5,.5}\pm 5.5}}$ $71.3_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $79.8_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $62.4$
VIM $77.2_{{\color[rgb]{.5,.5,.5}\pm 3.0}}$ $52.7_{{\color[rgb]{.5,.5,.5}\pm 3.8}}$ $63.9_{{\color[rgb]{.5,.5,.5}\pm 3.8}}$ $71.8_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $64.8_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $66.1$
@@
KNN $93.9_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $58.8_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $92.1_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $73.3_{{\color[rgb]{.5,.5,.5}\pm 0.6}}$ $70.3_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $77.7$
@@
RMDS $95.0_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $99.0_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $97.2_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $85.8_{{\color[rgb]{.5,.5,.5}\pm 0.7}}$ $89.2_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $93.2$
@@
RMDS $94.2_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $99.1_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $96.4_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $83.5_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $88.0_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $92.2$
@@
VIM $79.6_{{\color[rgb]{.5,.5,.5}\pm 2.5}}$ $50.4_{{\color[rgb]{.5,.5,.5}\pm 3.1}}$ $60.7_{{\color[rgb]{.5,.5,.5}\pm 1.7}}$ $78.6_{{\color[rgb]{.5,.5,.5}\pm 0.6}}$ $77.4_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $69.3$
@@
KNN $98.6_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $54.5_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $91.1_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $79.7_{{\color[rgb]{.5,.5,.5}\pm 0.6}}$ $77.4_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $80.3$
@@
SPROD $\textbf{98.8}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $\textbf{61.6}_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $\textbf{97.4}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $\textbf{82.1}_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $\textbf{85.3}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $85.0$
FPR@95$\downarrow$
@@
MDS $49.2_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $96.0_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $39.0_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $93.0_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $90.5_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $73.5$
RMDS $91.7_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $99.6_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $95.3_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $83.4_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $88.1_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $91.6$
@@
SPROD $\textbf{4.7}_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $\textbf{93.7}_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $\textbf{19.0}_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $\textbf{70.1}_{{\color[rgb]{.5,.5,.5}\pm 1.2}}$ $58.0_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ 49.1
@@
Energy $69.0_{{\color[rgb]{.5,.5,.5}\pm 9.8}}$ $40.3_{{\color[rgb]{.5,.5,.5}\pm 10.5}}$ $48.9_{{\color[rgb]{.5,.5,.5}\pm 5.0}}$ $85.4_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $85.6_{{\color[rgb]{.5,.5,.5}\pm 0.7}}$ $65.8$
@@
RMDS $91.2_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $99.8_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $96.8_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $71.0_{{\color[rgb]{.5,.5,.5}\pm 1.7}}$ $60.1_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $83.8$
@@
KNN $58.6_{{\color[rgb]{.5,.5,.5}\pm 2.0}}$ $\textbf{95.3}_{{\color[rgb]{.5,.5,.5}\pm 0.7}}$ $57.6_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $76.0_{{\color[rgb]{.5,.5,.5}\pm 1.7}}$ $71.9_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $71.9$
SHE $44.0_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $98.3_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $38.1_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $92.1_{{\color[rgb]{.5,.5,.5}\pm 1.2}}$ $62.1_{{\color[rgb]{.5,.5,.5}\pm 1.3}}$ $66.9$
@@
Energy $66.8_{{\color[rgb]{.5,.5,.5}\pm 2.3}}$ $43.1_{{\color[rgb]{.5,.5,.5}\pm 9.4}}$ $44.2_{{\color[rgb]{.5,.5,.5}\pm 7.0}}$ $\textbf{87.6}_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $81.6_{{\color[rgb]{.5,.5,.5}\pm 0.6}}$ $64.7$
MLS $67.0_{{\color[rgb]{.5,.5,.5}\pm 2.1}}$ $42.9_{{\color[rgb]{.5,.5,.5}\pm 9.3}}$ $44.4_{{\color[rgb]{.5,.5,.5}\pm 6.3}}$ $87.4_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $82.2_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $64.8$

### v2

@@ Appendix E Performance on Non-Spurious OOD Datasets
GNorm $90.7_{{\color[rgb]{.5,.5,.5}\pm 0.6}}$ $70.8_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $64.1_{{\color[rgb]{.5,.5,.5}\pm 2.1}}$ $7.3_{{\color[rgb]{.5,.5,.5}\pm 0.1}} 99$$9.0_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $84.4$
ReAct $86.8_{{\color[rgb]{.5,.5,.5}\pm 5.2}}$ $66.1_{{\color[rgb]{.5,.5,.5}\pm 7.5}}$ $50.0_{{\color[rgb]{.5,.5,.5}\pm 8.7}}$ $6.3_{{\color[rgb]{.5,.5,.5}\pm 0.1}} 99$$9.0_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $79.6$
VIM $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $9.6_{{\color[rgb]{.5,.5,.5}\pm 0.4}} 99$$9_{{\color[rgb]{.5,.5,.5}\pm 0.0}} 99.$$6_{{\color[rgb]{.5,.5,.5}\pm 0.0}} 99.$899.8
MDS $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $99.4_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $7.5_{{\color[rgb]{.5,.5,.5}\pm 0.0}} 99$.499.4
@@
SHE $96.4_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $8.6_{{\color[rgb]{.5,.5,.5}\pm 0.2}} 99$$9.1_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $94.3_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $98.5_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $97.4$
@@
GNorm $97.2_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $49.7_{{\color[rgb]{.5,.5,.5}\pm 1.1}}$ $59.6_{{\color[rgb]{.5,.5,.5}\pm 2.0}}$ $8.4_{{\color[rgb]{.5,.5,.5}\pm 0.1}} 99$$9.0_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $80.8$
ReAct $96.1_{{\color[rgb]{.5,.5,.5}\pm 1.5}}$ $53.9_{{\color[rgb]{.5,.5,.5}\pm 4.6}}$ $40.6_{{\color[rgb]{.5,.5,.5}\pm 5.9}}$ $8.0_{{\color[rgb]{.5,.5,.5}\pm 0.0}} 99$$9.0_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $77.5$
VIM $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $99.2_{{\color[rgb]{.5,.5,.5}\pm 0.7}}$ $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $9.7_{{\color[rgb]{.5,.5,.5}\pm 0.0}} 99$.899.8
MDS $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $99.7_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $8.0_{{\color[rgb]{.5,.5,.5}\pm 0.0}} 99$.599.5
@@
ReAct $57.9_{{\color[rgb]{.5,.5,.5}\pm 15.9}}$ $78.6_{{\color[rgb]{.5,.5,.5}\pm 7.4}}$ $65.0_{{\color[rgb]{.5,.5,.5}\pm 8.0}}$ $1.8_{{\color[rgb]{.5,.5,.5}\pm 0.2}} 99$$9.0_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $78.5$
VIM $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $9.8_{{\color[rgb]{.5,.5,.5}\pm 0.2}} 99$$9_{{\color[rgb]{.5,.5,.5}\pm 0.0}} 99.$$2_{{\color[rgb]{.5,.5,.5}\pm 0.0}} 99.$899.8
MDS $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $\textbf{100.0}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $98.8_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $6.5_{{\color[rgb]{.5,.5,.5}\pm 0.0}} 99$.199.1
@@
SHE $1.6_{{\color[rgb]{.5,.5,.5}\pm 0.4}} 99$$4_{{\color[rgb]{.5,.5,.5}\pm 0.1}} 99.$$9.5_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $97.0_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $98.3_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $97.2$
@@
We evaluate the performance of post-hoc methods in the non-spurious OOD (NSP-OOD) detection setting. For this evaluation, the OOD samples are drawn from a collection of standard benchmark datasets: SVHN [78], LSUN (resized)[79], and iSUN[80]. The ID context for each evaluation varies across the five datasets used in our primary spurious correlation experiments: Waterbirds (WB), CelebA (CA), UrbanCars (UC), Animals MetaCoCo (AMC), and Spurious ImageNet (SpI). Specifically, for each of these five ID contexts, the pretrained ResNet-50 backbone is employed, and the post-hoc OOD detection methods are subsequently set up using the features derived from this backbone.
The performance metrics reported in Tables 9 and 10 are the average performance when distinguishing each ID dataset from the NSP-OOD samples, averaged across SVHN, LSUN, and iSUN. This setup assesses the general OOD detection capability of the post-hoc methods when the primary challenge is not spurious correlations shared between ID and OOD, but the ID datasets still contain inherent biases. As observed in the tables, distance-based methods such as SPROD and KNN consistently achieve near-perfect scores across all metrics (AUROC, FPR@95, AUPR-In, and AUPR-Out) and ID contexts. Other methods like VIM and MDS also show strong performance.
Appendix F Ablation Study on SPROD Stages
@@
Appendix G Backbone Experiments
@@
We analyze the performance of each method using four complementary metrics (AUROC, FPR@95, AUPR-IN, and AUPR-OUT) to evaluate different aspects of OOD detection performance. Tables 11, 12, 13, and 14 summarize the performance of our post-hoc method across various backbone architectures, averaged over the five SP-OOD datasets studied in this work. Detailed results for each individual backbone across the different datasets are presented in the subsequent tables.
@@
Overall, we observe that SPROD consistently achieves strong performance across all backbone architectures, often with a notable margin. KNN emerges as the second-best method, suggesting that simpler approaches can be counterintuitively effective; however, its performance is highly sensitive to the choice of hyperparameters. MDS is also among the top-performing methods; it is a metric-based approach similar to SPROD but employs a more complex model by estimating class-specific covariance matrices. This added complexity, while potentially beneficial, may increase the risk of overfitting, especially in OOD settings. In addition, output-based methods show notable drops in performance on several benchmarks, highlighting their susceptibility to distributional shifts. Notably, SPROD is the only method that maintains stable performance across all evaluated settings, without experiencing major degradation under any backbone or dataset configuration. Among the evaluated backbones, ConvNeXt-B and Swin-B stand out as frozen feature extractors with superior performance for post-hoc OOD detection.
@@
MSP $68.1_{{\color[rgb]{.5,.5,.5}\pm 0.7}}$ $47.6_{{\color[rgb]{.5,.5,.5}\pm 2.9}}$ $40.2_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $2.4_{{\color[rgb]{.5,.5,.5}\pm 0.6}} 77$$7.3_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $61.1$
@@
ReAct $75.9_{{\color[rgb]{.5,.5,.5}\pm 4.1}}$ $0.8_{{\color[rgb]{.5,.5,.5}\pm 3.7}} 44$$4.0_{{\color[rgb]{.5,.5,.5}\pm 5.5}}$ $71.3_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $79.8_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $62.4$
VIM $77.2_{{\color[rgb]{.5,.5,.5}\pm 3.0}}$ $52.7_{{\color[rgb]{.5,.5,.5}\pm 3.8}}$ $63.9_{{\color[rgb]{.5,.5,.5}\pm 3.8}}$ $71.8_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $4.8_{{\color[rgb]{.5,.5,.5}\pm 0.2}} 66$.166.1
@@
KNN $93.9_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $58.8_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $92.1_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $73.3_{{\color[rgb]{.5,.5,.5}\pm 0.6}}$ $0.3_{{\color[rgb]{.5,.5,.5}\pm 0.1}} 77$.777.7
@@
RMDS $5.0_{{\color[rgb]{.5,.5,.5}\pm 0.0}} 99$$9.0_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $97.2_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $85.8_{{\color[rgb]{.5,.5,.5}\pm 0.7}}$ $89.2_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $93.2$
@@
RMDS $4.2_{{\color[rgb]{.5,.5,.5}\pm 0.1}} 99$$9.1_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $96.4_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $3.5_{{\color[rgb]{.5,.5,.5}\pm 0.9}} 88$$8.0_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $92.2$
@@
VIM $79.6_{{\color[rgb]{.5,.5,.5}\pm 2.5}}$ $50.4_{{\color[rgb]{.5,.5,.5}\pm 3.1}}$ $60.7_{{\color[rgb]{.5,.5,.5}\pm 1.7}}$ $8.6_{{\color[rgb]{.5,.5,.5}\pm 0.6}} 77$$7.4_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $69.3$
@@
KNN $98.6_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $54.5_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $91.1_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $9.7_{{\color[rgb]{.5,.5,.5}\pm 0.6}} 77$$7.4_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $80.3$
@@
SPROD $\textbf{98.8}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $\textbf{61.6}_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $\textbf{97.4}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $\textbf{82.4}_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $\textbf{85.3}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $85.1$
[…]
@@
MDS $49.2_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $96.0_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $_{{\color[rgb]{.5,.5,.5}\pm 0.3}} 93.0±0.39$$_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $90.5_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $73.5$
RMDS $1.7_{{\color[rgb]{.5,.5,.5}\pm 0.2}} 99$$9.6_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $95.3_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $3.4_{{\color[rgb]{.5,.5,.5}\pm 0.9}} 88$$8.1_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $91.6$
@@
SPROD $\textbf{4.7}_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $\textbf{93.7}_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $\textbf{19.0}_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $\textbf{69.5}_{{\color[rgb]{.5,.5,.5}\pm 1.2}}$ $58.0_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ 49.0
@@
Energy $69.0_{{\color[rgb]{.5,.5,.5}\pm 9.8}}$ $40.3_{{\color[rgb]{.5,.5,.5}\pm 10.5}}$ $48.9_{{\color[rgb]{.5,.5,.5}\pm 5.0}}$ $85.4_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $_{{\color[rgb]{.5,.5,.5}\pm 0.7}} 65.865.8$
@@
RMDS $1.2_{{\color[rgb]{.5,.5,.5}\pm 0.1}} 99$$9.8_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $96.8_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $71.0_{{\color[rgb]{.5,.5,.5}\pm 1.7}}$ $60.1_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $83.8$
@@
KNN $58.6_{{\color[rgb]{.5,.5,.5}\pm 2.0}}$ $\textbf{95.3}_{{\color[rgb]{.5,.5,.5}\pm 0.7}}$ $57.6_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $76.0_{{\color[rgb]{.5,.5,.5}\pm 1.7}}$ $_{{\color[rgb]{.5,.5,.5}\pm 0.1}} 71.971.9$
SHE $44.0_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $98.3_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $38.1_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $92.1_{{\color[rgb]{.5,.5,.5}\pm 1.2}}$ $2.1_{{\color[rgb]{.5,.5,.5}\pm 1.3}} 66$.966.9
@@
Energy $66.8_{{\color[rgb]{.5,.5,.5}\pm 2.3}}$ $3.1_{{\color[rgb]{.5,.5,.5}\pm 9.4}} 44$$4.2_{{\color[rgb]{.5,.5,.5}\pm 7.0}}$ $\textbf{87.6}_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $81.6_{{\color[rgb]{.5,.5,.5}\pm 0.6}}$ $64.7$
MLS $67.0_{{\color[rgb]{.5,.5,.5}\pm 2.1}}$ $2.9_{{\color[rgb]{.5,.5,.5}\pm 9.3}} 44$$4.4_{{\color[rgb]{.5,.5,.5}\pm 6.3}}$ $87.4_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $82.2_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $64.8$